### v1

# Towards Operational Automated Greenhouse Gas Plume Detection

Abstract
Operational deployment of a fully automated greenhouse gas (GHG) plume detection system remains an elusive goal for imaging spectroscopy missions, despite recent advances in deep learning approaches. With the dramatic increase in data availability, however, automation continues to increase in importance for natural and anthropogenic emissions monitoring. This work reviews and addresses several key obstacles in the field: data and label quality control, prevention of spatiotemporal biases, and correctly aligned modeling objectives. We demonstrate through rigorous experiments using multicampaign data from airborne and spaceborne instruments that convolutional neural networks (CNNs) are able to achieve operational detection performance when these obstacles are alleviated. We demonstrate that a multitask model that learns both instance detection and pixelwise segmentation simultaneously can successfully lead towards an operational pathway. We evaluate the model’s plume detectability across emission source types and regions, identifying thresholds for operational deployment. Finally, we provide analysis-ready data, models, and source code for reproducibility, and work to define a set of best practices and validation standards to facilitate future contributions to the field.
keywords:
Plume Detection , Greenhouse Gas , Machine Learning , Imaging Spectroscopy[][nocite]build/opsghgml_supplemental
[jpl]organization=Jet Propulsion Laboratory, California Institute of Technology, city=Pasadena, state=CA, postcode=91101, country=USA
[cm]organization=Carbon Mapper Inc., city=Pasadena, state=CA, postcode=91101, country=USA
[pri]organization=Princeton University, city=Princeton, state=NJ, postcode=08544, country=USA
[cit]organization=California Institute of Technology, city=Pasadena, state=CA, postcode=91125, country=USA
[col]organization=Columbia University, city=New York, state=NY, postcode=10027, country=USA
1 Introduction
Modern airborne & spaceborne imaging spectrometers such as the Airborne Visible-Infrared Imaging Spectrometer-Next Generation (AVIRIS-NG) (Hamlin et al., 2011; Chapman et al., 2019), the Global Airborne Observatory (GAO) (Asner, 2007; Asner et al., 2012), the Earth surface Mineral dust source InvesTigation (EMIT) (Green et al., 2020), and Carbon Mapper Coalition’s Tanager-1 (Zandbergen et al., 2023) capture finely-sampled observations with spectral resolution sufficient to detect GreenHouse Gas (GHG) emissions with discriminative absorptions in the Visible to ShortWave InfraRed (VSWIR) and spatial resolution sufficient to resolve meter-scale surface infrastructure elements.
While current sensor platforms equipped with modern GHG retrieval approaches yield products that are valuable for plume identification and site-specific emissions quantification tasks, they are also prone to generating false enhancements. The presence of these false enhancements has to date prevented the operational deployment of fully automated plume detection and localization tasks.
Spatially-informed supervised machine learning models such as Convolutional Neural Networks (CNNs) provide a potential solution to automatically detect, localize, and delineate plumes from GHG retrievals. CNN-based approaches have proven successful in a variety of complex image classification, object detection, and pixelwise segmentation benchmark data sets (Russakovsky et al., 2014; Everingham et al., 2010; Krizhevsky, 2009). Recent studies have claimed impressive results using CNN-based plume detection systems driven by GHG image products derived from data captured by airborne and spaceborne instruments (Joyce et al., 2022; Schuit et al., 2023; Bruno et al., 2023; Vaughan et al., 2023; Kumar et al., 2023; Rouet-Leduc and Hulbert, 2024).
Advancing ML-driven plume detection capabilities to global scales demands informed optimization and validation strategies to ensure deployed models generalize operationally. Assessing the generalization capabilities of a supervised ML model involves distinguishing settings where we can trust the outputs produced by the model from those where we do not. While this is indeed a fundamentally difficult task, evaluating the capabilities of a proposed plume detector relative to its associated training and validation data with respect to the following criteria has proven instructive in our work:
1.
Do the validation procedures and performance metrics accurately and unambiguously capture the performance of the model on the available data?
2.
Are the available data sufficiently representative of the diversity of observations we expect to encounter when deployed operationally?
We demonstrate in this work that effectively constructing and deploying an operational plume detector is an inherently interdisciplinary task, and that collaborative analysis performed by both ML practitioners and domain experts is crucial to operational plume detection capabilities. We show that systematic, informed analysis of model predictions, their associated GHG retrieval products and the settings in which they were captured can effectively reveal unexpected model or data-driven biases that impact generalization performance and methodological issues that corrupt training and validation procedures that produce misleading or uninterpretable results.
We address these issues in this work and in our previous efforts Bue et al. (2023, 2024); Lee et al. (2021, 2022, 2023) by guided by the following research questions:
1.
What science objectives must an automated plume detection system address to be operationally effective?
2.
How should we frame plume detection as a machine learning problem to best progress towards achieving operational objectives?
3.
How do we rigorously and unambiguously characterize the capabilities of a plume detection system driven by machine learning?
We aim to provide a basis towards advancing operational capabilities and identifying challenges involved in applying machine learning methods for automated plume detection. By collectively addressing these issues we hope to achieve more rigorous validation leading to an operational system for this critical climate problem.
This paper is organized as follows: Section 2 provides background information on the CMF retrieval product, GHG plumes and their labels, and definitions of the plume detection task. Section 3 describes the plume detection system, including quality control, sampling, and validation. Section 4 describes this work’s data sets and models, and Section 5 presents the corresponding plume detection results. Finally, Section 7 puts these results in context of limitations in existing literature, and we conclude in Section 8.
2 Background
To define operational requirements for a plume detection system, we first must understand the data and labels that will be used to construct and validate the system. Equipped with this knowledge, we can effectively construct protocols to collect representative samples to train and test the model, define optimization routines that account for task-specific sources of bias and uncertainty, and select capable validation metrics that provide realistic estimates of generalization performance. While the data and results described in this work are focused on matched filter-based retrievals of methane derived from imaging spectrometer observations, the challenges we describe are generally relevant to similar plume detection systems, including those based on image data captured by analogous remote sensing platforms such as PRISMA and EnMAP (Guanter et al., 2023; Roger et al., 2024), targeting other greenhouse gases such as or , and retrievals derived using alternative algorithms such as mag1c (Foote et al., 2020).
2.1 GHG Retrieval Image Products
We derive pixelwise GHG concentrations from calibrated imaging spectrometer radiance observations using the Columnwise Matched Filter (CMF) (Thompson et al., 2015) retrieval approach. Based largely on the methods developed and refined by scientists at Los Alamos National Labs starting in the late 1990s (Villeneuve et al., 1999; Theiler, 2012), the CMF is conceptually simple, computationally efficient, and produces retrievals that are both numerically stable across repeat observations of the same sites and emission sources and consistent with coincident measurements by complementary sensors (e.g., comparisons with dynamic aviation soundings collected during Four Corners campaign (Thompson et al., 2015)) and with measurements from controlled release experiments (Thorpe et al., 2017, 2023, 2016; Duren et al., 2019; El Abbadi et al., 2024; Sherwin et al., 2023; Ayasse et al., 2023). The CMF retrieval computes the pixelwise enhancement of VSWIR absorption features associated with a MODTRAN-simulated GHG transmittance spectrum in units relative to multivariate Gaussian backgrounds, each estimated columnwise from the observations of the independent detector elements in the sensor’s focal plane array. CMF products derived from AVIRIS-NG and GAO data are highly analogous—both are based on the approach described in Thompson et al. (2015) that retrieves pixelwise concentrations according to the depth of the SWIR2 absorption feature of airborne spectrometer radiance observations captured at GSD [1,10] m. Any differences between the instruments are dwarfed by atmospheric, geometric, and radiative transfer assumptions within the retrieval. EMIT observations are observed at 60 m GSD and EMIT CMF products are derived using the same matched-filter-based algorithm used with the airborne data. The operational EMIT L2 product uses all VSWIR wavelengths rather than just the SWIR2, is more stable due to its environment in orbit, and leverages a Look-Up Table (LUT) to select an appropriate target spectrum for the retrieval that accounts for varying concentration levels and solar zenith angles (Foote et al., 2021; Thorpe et al., 2023). Filtering of cloud, water, and thermal anomalies from the calculation of the covariance also avoids artifacts in the retrieval.
2.2 GHG Plume Types
GHG plumes manifest as spatially contiguous clusters of pixels distinguishable by higher concentration levels than their surrounding background pixels. Figure 1 shows several example plumes from distinct source types observed in airborne imaging campaigns. The type of emission source and the emission rate, adjusted by dispersion from local winds, are the primary factors determining the concentration level and morphology of a given plume. Although distinct infrastructure types from various emission sources produce plumes with significant variability, for the purposes of detection we can broadly group source emitters into two categories based on the characteristics of the plumes they emit.
GHG plumes from oil & natural gas infrastructure and industrial facilities tend to produce “point source” plumes with localized, concentrated emissions that can be traced back to a single specific location. The shape of point source plumes is largely governed by local wind conditions. In contrast, “diffuse sources” from agricultural infrastructure (e.g., manure, enteric fermentation, rice cultivation) and solid waste management facilities (e.g., landfills) produce plumes representing the collective emissions of multiple sources in close proximity of one another, or plumes emitted by area sources lacking distinct points of origin.
2.3 Plume Labels
This work leverages plumes identified by domain experts as described in Suppl. Sections LABEL:apx:qc and LABEL:apx:airborne. Constructing and validating a ML-driven GHG plume detector also requires pixelwise labels to teach the model to distinguish plume pixels from background pixels. However, plumes are not discrete objects with well-defined boundaries, and manually labeling individual plume pixels is a difficult and subjective task, particularly near plume boundaries. One alternative to requiring users to manually label individual plume pixels is to use the CMF enhancements to deterministically generate “CMF-guided” plume labels consisting of highly concentrated pixels near the origin point of each user identified plume. While the CMF-guided labels may be less subjective than manually defined labels, their quality can vary significantly by scene according to the local contrast in values between plumes and background enhancements or the presence of false enhancements adjacent to real plumes, and consequently require comprehensive manual inspection to refine or reject uninformative instances.
Figure 2 shows two example plume candidates (left column) with corresponding labels provided by four independent experts (center column), along with the deterministic CMF-guided plume labels for each plume according to increasing thresholds (right column). All four users agree that each enhancement represents a real plume, but the pixelwise masks they each provide disagree with respect to the boundaries where plume pixels end and background pixels begin (center column). In contrast, the CMF-guided labels are more conservative, capturing only the highly concentrated pixels near the origin point associated with each plume.
2.4 Background and False Enhancements
Even in targeted GHG monitoring campaigns involving repeated measurements of known emission sources, pixels representing GHG plumes are extremely rare. For instance, labeled plume pixels account for less than 1% of the science pixels captured in the CMFs from the 2018-2020 airborne AVIRIS-NG and GAO imaging campaigns. The remaining pixels represent enhancements where the target GHG is absent or concentrated at a level too low to be detectable by the retrieval algorithm or measurable by the sensor. These enhancements generally fall into one of two categories: Background Enhancements and False Enhancements, typified by the spatial distribution of their enhancements, how often they occur in each scene, and the factors (real or artificial) that produce them.
Background Enhancements: The majority of pixels in a typical scene are common background enhancements representing retrievals of surface materials lacking characteristic GHG absorption features varying with scenewise imaging geometry, albedo, illumination conditions, and atmospheric state. Most background enhancements exhibit small values that fall within the range where GHG concentrations are too low to distinguish them from instrument noise (e.g., 250 in typical airborne CMFs). By definition, the CMF column concentration should follow a zero-mean distribution, with the diffuse background emerging as a consequence of instrument noise characteristics and environmental conditions. Though higher concentration retrievals occasionally occur in regions dominated by common background enhancements, they tend to be spatially sparse. Consequently, regions consisting primarily of common background enhancements consist of mostly low concentration pixels with sporadic high enhancement pixels that lack cohesive spatial structure, making such regions easily distinguishable from real GHG plumes.
False Enhancements: False enhancements are a nuisance class of background enhancements with concentration levels within the range of real GHG plumes (exceeding 500-1000) caused by factors unrelated to the presence of GHG emissions. One common driver involves confusers, including surface spectroscopy that mimics methane absorption features at the spectral sampling interval of the instrument and anomalous observations, such as high albedo materials and thermal anomalies. Another involves artifacts caused by limitations of the instrument or retrieval method, such as corrupt background covariance estimates due to flares and sample size limitations from short flightlines. Although false enhancements are relatively rare in terms of pixelwise coverage (we estimate false enhancements represent 1-5% of the retrieved pixels in a typical CMF scene) false enhancements that manifest at spatial scales and with spatial structure similar to real plumes constitute the majority of false positive plume detections observed in this work, and can be difficult even for domain experts to distinguish from real GHG plumes. False enhancements can severely restrict plume detectability when they occur near real emission sources and merge with pixels representing observed plumes, thereby altering their spatial distribution and concentration levels. Additionally, in some rare, pathological cases, false enhancements can cover the spatial extent of an entire scene (see examples in Suppl. Section LABEL:apx:gao_penn and Suppl. Figure LABEL:fig:penn_tiles).
It is worth noting that, in cases of low sensitivity observations, it is also possible to identify false negative enhancements. While less problematic and even rarer than the false positive enhancements discussed above, it can cause challenging plume morphologies; for example, a plume blowing over a river may appear to be cut in half where methane was not retrieved due to the low albedo.
Figure 3 shows example background and false enhancements. Common background enhancements (columns 1 and 2) consist of low concentration pixels and sporadic, sparsely distributed high enhancement pixels. False enhancements from confuser materials or albedo variation (columns 3-4) are typically correlated with visible surface features, while false enhancements from artifacts (column 5) tend to be uncorrelated with surface features, typically following downtrack imaging geometry (e.g., columnwise retrieval artifacts) or with both downtrack & crosstrack imaging geometry simultaneously (e.g., Fourier ringing from thermal saturation in SWIR channels caused by oil & natural gas flares (Foote et al., 2018)).
2.5 Operational Plume Detection
An automated plume detection system is operationally effective for image products measuring emissions of a target GHG (e.g., CMF images of or ) observed by a target remote sensing platform (e.g., AVIRIS-NG, EMIT) if, when applied to arbitrary target images observed in nominal imaging conditions, the plume detection system performs comparably to domain experts with respect to the following objectives:
1.
Instance Detection: Detect and localize identifiable target plumes and reject identifiable false enhancements.
2.
Pixelwise Segmentation: Delineate pixelwise boundaries of detected plumes.
Both of these objectives are achievable using current GHG image retrieval approaches applied to data captured by modern airborne and spaceborne imaging spectrometer platforms. Prior manual plume identification efforts have shown that human experts can consistently identify, localize and delineate most plumes through visual inspection of the CMF retrieval images we consider in this work. Comprehensive manual analysis of retrieval images captured in airborne imaging campaigns has also informed improvements to GHG retrieval algorithms, suppressing common classes of false enhancements, and providing diagnostic capabilities to quantify retrieval sensitivity and uncertainty.
Requiring a plume detector to perform comparably to human experts makes for a challenging machine learning problem. However, it also provides constraints that simplify validation efforts. Specifically, we do not include scenes in model training where domain experts are unable to exhaustively identify all candidate plume enhancements in the scene with adequate confidence. We also require all plume candidates be identifiable using only the GHG retrieval images we use to train and validate our models. Ambiguous plume candidates that can only be resolved by cross-referencing external data sets (e.g., high spatial resolution basemap images) or prior measurements of the same site are excluded from our experiments. Evaluating pixelwise plume delineation accuracy when human experts disagree regarding which pixels distinguish plumes from background pixels is difficult to interpret, so we avoid it. Details regarding the quality control criteria we use to exclude low SnR scenes and unidentifiable plumes are provided in Section 4.1, and Supplemental Section LABEL:sub:angqc.
While a fully-operational system must be able to flag data quality issues such as low SnR scenes or ambiguous detections for manual review or reprocessing, relying on the plume detector for such tasks is often a waste of resources, as simple triage protocols to account for and handle problematic scenes could be more effective at identifying known issues. We provide a summary of several data triage routines in Supplemental Section LABEL:sub:angqc.
3 Methods
System Overview: We consider plume detection systems of the form shown in Figure 4 in this work. Provided a set of labeled scenes consisting of GHG retrieval image products and their associated pixelwise plume labels, our objective is to train a plume detector model to estimate the probability (or “plume salience”) that each pixel in the image represents a plume. Supervised machine learning models achieve this by learning to map labeled plume image pixels to high salience values and unlabeled “background” image pixels to low salience values. In this work, we focus on CNN-based plume detectors which learn a hierarchy of spatial filters that capture disscriminative features of labeled plumes and unlabeled background enhancements using fixed-size image tiles extracted from our labeled CMF scenes. Once trained, CNNs are capable of efficiently generating spatially-informed pixelwise predictions (or “plume salience maps”) based on CMF images with arbitrary input dimensions.
Quality Control: We rely on comprehensive quality control procedures (see Suppl. Sections LABEL:sub:angqc and LABEL:sub:emitqc) to ensure that our models are trained and validated on high quality, exhaustively labeled CMF image products whose constituent plume labels are readily identifiable and distinguishable from (unlabeled) background enhancements by domain experts without additional context.
Tilewise Training & Scenewise Validation: Once deployed, we assume an operational plume detector will be used to detect and delineate plumes in “full scene” CMF image products, and the performance of candidate plume detector models should be validated on full scene CMF products to replicate operational conditions. While the CNN models described in this work are functionally capable of generating full scene predictions, scenewise validation of plume predictions is not standard practice. This is due to the fact that CNNs are trained “tilewise” using collections of sub-image tiles with fixed dimensions extracted from full CMF scenes. Once trained, such models are typically also evaluated “tilewise” by applying the model to test tiles with the same image dimensions.
We apply our tilewise-trained models to generate salience maps of the subset of our test scenes that meet strict “scenewise” quality-control requirements (see Suppl. Section LABEL:sub:angqc for details regarding quality control criteria). For these scenes, we measure prediction accuracy on all valid pixels in each scene with respect to pixelwise plume segmentation and plume instance detection (see Suppl. Section LABEL:sub:instance for details regarding measuring plume instance detection accuracy).
To ensure our models and results are not spatiotemporally biased, we use the bounding boxes of the georeferenced CMF images to stratify our data into spatially disjoint training and test scenes. For airborne campaigns, we compute the graph of pairwise Intersection-over-Union (IoU) scores with respect to the scenes observed in each campaign. Each connected component of the resulting IoU graph identifies a group of overlapping flightlines that are spatially disjoint from all the remaining flightlines. For the spaceborne EMIT data, we group scenes according to their respective UTM & MGRS zone ids. We generate spatially disjoint, class stratified training and test partitions by assigning the labeled plume and background tiles associated with each group of flightlines to either the training set or the test set via the GroupShuffleSplit function provided in the scikit-learn package (Pedregosa et al., 2011). For each campaign, we designate roughly 75% of the flightlines (and their associated tiles) as training data, and the remaining 25% of the flightlines as test data.
Spatially-stratified Tile Sampling: Similar to the scene-wise stratification, we need to ensure that the sampled distribution of CMF tiles is well distributed and non-overlapping within a given scene. We use a spatially-stratified tile sampling strategy that prioritizes sampling background tiles in regions with high concentration levels, and incorporates controls that explicitly prevent inter-class tile overlap (i.e., plume and background tiles may never overlap) while permitting nominal within class overlap (i.e., background tiles may overlap other background tiles). Our procedure begins by extracting candidate plume tiles at the center of mass of each labeled plume ROI. We then exclude any pixelwise plume labels that lie within the current plume tile boundaries, which partitions any large plume ROIs into a new set of sub-plume ROIs. We repeat this process until we cannot sample any more spatially disjoint plume tiles centered on distinct sub-plume ROIs. After sampling all available plume tiles, we use the CMF-guided background labels to randomly sample background tiles that roughly cover the remainder of the scene. We ensure that background tiles do not overlap any existing plume tiles or any pixelwise plume labels outside of the selected plume tiles. However, we do allow background tiles to overlap other background tiles in the scene by at most 10%, and we also allow background tiles to overlap nodata regions (most commonly, scene boundaries) by at most 50%. We found that allowing some amount of background-background overlap allows for a more comprehensive sampling of the background enhancements within each scene and exposes the model to observations near scene boundaries that are inevitably present in full CMF images.
Evaluating Plume Detection and Segmentation Metrics In addition to tilewise performance metrics, we report plume detection and segmentation metrics calculated over entire scenes/flightlines. Pixelwise segmentation metrics measure how closely pixelwise predictions align with labeled plume boundaries, while instance-level detection metrics measure how many plume instances were successfully detected and how many false plume detections occurred. High pixelwise accuracy does not imply high instance detection accuracy (and vice-versa), so both instance-level plume detection and pixelwise segmentation metrics are essential to unambiguously assess plume detector performance. Pixelwise segmentation metrics are reported simply as precision, recall, and F1-Scores over all pixels, while instance-level detections are reported based on whether a detection mask overlaps a label mask; this is defined in detail in Suppl. Section LABEL:sub:instance.
4 Data & Models
4.1 Data
Table 5 provides a summary of all of the CMF data sets we consider in this work. We summarize relevant details regarding the data sets used for training and validating our airborne (AVIRIS-NG and GAO) and spaceborne (EMIT) plume detectors here. Detailed column definitions for Table 5 are provided in Section A, and quality control levels are described in Suppl. Section LABEL:apx:qc. The distribution of source sectors for plumes observed in airborne imaging campaigns are provided in Section B.
4.2 Airborne Plume Detector Data
AVIRIS-NG Multicampaign Data Set: Except when otherwise specified, all plume detection experiments described in this work involving airborne imaging campaigns are based on the curated, quality controlled set of labeled CMF images derived from data captured by the AVIRIS-NG airborne imaging spectrometer in the 2018 CalCH4 campaign Duren et al. (2019), 2019 Permian Campaign Cusworth et al. (2021), and 2020 COVID Thorpe et al. (2023) imaging campaigns.
GAO Multicampaign Data Set: We evaluate the generalization capabilities of the airborne plume detectors on test data consisting of labeled CMF images derived from data captured by the GAO airborne imaging spectrometer in the 2019 Permian, 2020 CARB, and 2021 Denver imaging campaigns (Cusworth et al., 2022).
Unique Campaigns: CMF image products derived from data captured in the 2015 AVIRIS-NG Four Corners campaign and the 2021 GAO Penn campaign exhibit characteristics that are distinct from CMF images derived from data captured in other airborne campaigns. We summarize how these distinctions impact plume detection capabilities and the interpretation of plume detection result in Suppl. Section LABEL:apx:bias_region, but we exclude the observations from these anomalous campaigns from all of our remaining experiments.
4.3 EMIT Plume Detector Data
EMIT VISIONS Data Set: All plume detection experiments involving EMIT observations were trained and tested on observations from the EMIT VISIONS Data Set. The EMIT VISIONS data set consists of the expert curated, high SnR CMF image products derived from data captured by the spaceborne imaging spectrometer EMIT between 2022-2024 that contain high-confidence methane plumes delivered to the public-facing EMIT VISIONS portal 111https://earth.jpl.nasa.gov/emit/data/data-portal/Greenhouse-Gases/ that were formally-reviewed & approved by multiple reviewers using the Multi-Mission Geographic Information System (Soliman et al., 2025).
4.4 Models
CNN Architectures
Our experiments leverage two widely-used and well-established CNN architectures: GoogLeNet and U-Net. Model architecture, training procedures, and loss functions referenced are described in full in Suppl. Section LABEL:apx:ml.
1.
GoogLeNet: The GoogLeNet architecture (Szegedy et al., 2015) incorporates several “inception blocks” designed to learn convolutional filters capturing discriminative image patterns at multiple spatial scales. Designed for image classification tasks, we train GoogLeNet using tiles extracted from CMF image scenes with tilewise plume or background labels determined by the presence of plume pixels in each tile. We compare the capabilities of GoogLeNet-based plume image classifier models to those of the recently proposed MethaNet (Jongaramrungruang et al., 2022) plume image classifier model on a variety of plume image classification experiments (see Sections LABEL:apx:bias_region and LABEL:apx:bias_distrib). To perform pixelwise plume segmentation, we convert our trained GoogLeNet image classifier to a “shift and stitch” Fully Convolutional Network (FCN) via the procedure described in Shelhamer et al. (2015). This allows us to efficiently apply GoogLeNet image classifiers to full-scene, pixelwise plume segmentation and instance detection tasks. We also apply model anti-aliasing as described in Zhang (2019) to avoid shift invariance artifacts in produced masks (Lee et al., 2020).
2.
U-Net: The U-net architecture Ronneberger et al. (2015) is among the most widely-used CNN architectures to date. Originally designed for medical applications involving pixelwise ”semantic segmentation” tasks Siddique et al. (2021), U-net is the base plume detector model used in several recent papers (Bruno et al., 2023; Růžička et al., 2023). The U-net architecture learns detailed image patterns directly from pixelwise labeled training images by connecting each convolution/downsampling layer to a matching deconvolution/upsampling layer. These ‘skip connections” allow the model to learn discriminative image filters at multiple spatial scales (via convolution & pooling) while simultaneously learning to generate accurate pixelwise predictions at native spatial resolution by combining the outputs of learned multiscale filters.
Scenewise Plume Detector Models
For our experiments involving airborne plume detection, we evaluated the capabilities of the following scenewise plume detector models with respect to plume instance detection and pixelwise segmentation objectives.
1.
Tilewise Model: “Shift & Stitch” GoogLeNet FCN. Since this model was trained for image classification tasks using tilewise (rather than pixelwise) labels, it offers advantages for plume instance detection but is ill-suited for pixelwise segmentation due to very coarse masks.
2.
Pixelwise Model: Standard U-net optimized with purely pixelwise objective. Generates detailed pixelwise predictions but is ill-suited for instance detection due to a large number of false positives.
3.
Multitask Model: Identical U-net model as pixelwise model optimized with multitask segmentation & classification objective. Provides a compromise between plume instance detection and pixelwise segmentation objectives.
Based on our airborne plume detection results with the models defined above, we trained and validated a separate multitask spaceborne plume detector model using the EMIT visions dataset and provide a comparative analysis of our airborne versus spaceborne plume detection/segmentation results.
5 Results
5.1 Scenewise Plume Instance Detection and Segmentation
Figure 7 shows two example plumes and the corresponding plume salience maps produced by the tilewise, pixelwise, and multitask models. The tilewise model produces detections that are typically much larger than the extent of the each plume, and tilewise detections occasionally merge plumes and neighboring false enhancements. However, the tilewise model generates fewer false positive instance detections than the pixelwise and multitask models, and also comes with the distinct advantage of not requiring pixelwise labels during training.
The pixelwise model generates detections that fit the boundaries of high concentration plume enhancements well, but also generates more spurious false detections than the other models. Most of these false detections correspond to small, ambiguous enhancements or image artifacts that are more effectively rejected by the other two models. However, the pixelwise model is more effective than the other models for plume bodies split into several adjacent disconnected components.
The multitask model provides a compromise between instance detection and pixelwise segmentation, and generates compact plume detections that fit high contrast plume boundaries more accurately than the other models, while also generating fewer false positive instance detections than the pixelwise model. However, because the multitask model generates more conservative detections than the other models, these improvements come at the cost of a higher false negative rate.
Comparing AVIRIS-NG Plume Detector Models: While the tilewise-trained model yields the highest instance detection accuracy overall, it performs poorly pixelwise due to the coarse segments it generates. Consequently, excluding small/weak detection candidates produces the smallest increases in both pixelwise and instance-level accuracy of the three models.
The pixelwise-trained model gives a slight improvement in pixelwise accuracy, but produces the worst overall instance detection accuracy due to the quantity of small false positive detections it generates. Consequently, excluding ambiguous (small/weak) detection candidates produced by the pixelwise model provides the most substantial increase in detection accuracy of the three models.
The multitask model is less prone to overfitting to ambiguous/uncertain plume boundary pixels in the training data, which accounts for its increase in pixelwise accuracy over the pixelwise-trained model. It also performs competitively with the tilewise model in terms of instance detection, particularly when we exclude small/weak ambiguous detections.
Comparing AVIRIS-NG and EMIT Results: With respect to pixelwise segmentation performance, AVIRIS-NG pixelwise mispredictions are dominated by false positives (precision recall) while false negatives dominate the EMIT pixelwise mispredictions (precision recall). Several factors contribute to the mismatched pixelwise performance of the AVIRIS-NG and EMIT models on their respective validation scenes.
The primary difference lies in the methodology used to define pixelwise plume labels for the airborne platforms versus the EMIT platform; EMIT plume labels are inclusive by design, while AVIRIS-NG plume labels are intentionally conservative. EMIT plume labelers are instructed to include a small pixel buffer around any plume they annotate, so the labels for each EMIT plume instance will capture some background pixels outside of each plume’s boundary. These low-concentration pixels account for the higher pixelwise false negative rate (lower recall) on the EMIT plumes. In contrast, the cmf-guided AVIRIS-NG plume labels only include high-concentration pixels near each plume’s origin point that exceed our 500 background enhancement threshold. As a result, the pixelwise labels of AVIRIS-NG plume instances often do not capture the full spatial extent of each plume, often excluding boundary pixels and “disconnected” plume components. These unlabeled pixels account for the higher pixelwise false positive rate (lower precision) on the AVIRIS-NG validation scenes. Additionally, the EMIT VISIONS data set only includes high SnR CMFs approved by multiple reviewers, while the AVIRIS-NG validation scenes were inspected and reviewed by fewer reviewers overall.
While differences in plume labeling methodologies broadly account for the mismatched distributions of pixelwise prediction errors between AVIRIS-NG and EMIT results, sensor and retrieval-specific differences help explain the gap in pixelwise precision between the AVIRIS-NG and EMIT results. EMIT is a more efficient sensor than AVIRIS-NG, and CMFs derived from EMIT data use an improved, full-VSWIR-based retrieval compared to the original SWIR2-only retrieval approach applied to AVIRIS-NG data from early airborne imaging campaigns. Compared to the AVIRIS-NG CMF products, plume and background pixels tend to be better separated in terms of local pixelwise contrast, and fewer false enhancements are present in EMIT CMFs. This is also reflected in the EMIT instance detection performance, where the multitask model performs better in both precision and recall.
5.2 Evaluating Regional Generalization
We evaluate generalization performance across distinct regions observed in AVIRIS-NG and GAO airborne imaging campaigns. However, since we do not have scenewise quality-controlled data for the GAO campaigns, we restrict our evaluation to tilewise instance classification performance results based on the Tilewise-trained model in this section.
Figure 8 summarizes the tilewise prediction accuracy of the AVIRIS-NG multicampaign model applied to the AVIRIS-NG training/test tiles and the GAO test tiles detailed in Table 5. In all cases the FPR dominates—typically by a factor of 2-3 the FNR, indicating that false positives constitute most of the prediction errors.
Figure 9 shows example false positive mispredictions produced by the AVIRIS-NG multicampaign model. All are false enhancements with concentration levels consistent with real plumes, though their morphologies vary according to the the phenomena that produce them. For instance, linear false enhancements that follow the downtrack imaging direction are MF artifacts produced when outliers corrupt the background covariance matrix used by the retrieval algorithm, while other linear false enhancements are caused by confuser materials (e.g., agricultural plastics, fertilizer, rooftop, or road surfaces with absorptions consistent with some hydrocarbons) coupled with sharp local changes in albedo.
We note that our comprehensive tile sampling strategy which selects tiles that cover the spatial extent of each scene plays a crucial role to ensure validation metrics capture realistic estimates of generalization performance. Due to the rarity of false enhancements, without comprehensive tile sampling, such phenomena would not be discovered.
Investigating the characteristics of plume mispredictions (false negatives) also yields informative diagnostics to assess the capabilities of the plume detector. Figure 10 shows randomly-selected false negatives from the AVIRIS-NG Four Corners, Cal and Permian campaigns, and the GAO Permian and CARB campaigns. Some mispredictions are due to small (A.3, B.4-B.5) or indistinct (A.4-5,B.2-3) plumes or spatially diffuse plumes surrounded by comparably high background enhancements (A.1, A.2). These likely approach the detection limits of the current plume detector model or retrieval algorithm. Others are likely driven by the presence of adjacent surface confusers (A.4-5, B.4-6) and/or adjacent columnwise artifacts (B.2) that the plume detector is trained to reject.
6 Measuring Airborne Plume Detectability
Having shown that the multicampaign model yields satisfactory results when applied to observations from similar regions, we now turn to the task of measuring plume detectability based on the characteristics of the labeled test plumes identified in those regions. As discussed in Section 2.5, we do not consider instances where enhancements fall below the range of instrument detectability. In the context of plumes observed by airborne spectrometers, our analysis below aims to address the following questions:
1.
How do plume size and concentration impact plume detectability?
2.
Can we reliably detect plumes generated by some classes of emission sources more reliably than others?
We base our assessment on the plumes observed in all airborne campaigns excluding those from the AVIRIS-NG Four Corners and GAO Penn campaigns. We exclude these due to their unique challenges and characteristics described in Suppl. Section LABEL:apx:bias_region.
Figure 11 gives the tilewise FNR scores for all AVIRIS-NG and GAO test plumes listed in Table 5 with respect to the average concentration of their constituent pixels (x-axis) versus the total area of each plume (y-axis). Most mispredictions occur at low average pixelwise concentrations () where plume and background enhancements are numerically similar. Weak plumes with small spatial footprints (area ) are the most commonly mispredicted (FNR=34.6%) since they are difficult to distinguish from false enhancements of similar size—such as the high albedo false enhancements shown in 9.
Figure 13 gives the average FNR (top) and the average number of false negatives (bottom) per-campaign with respect to the source emission sectors of the identified plumes in the selected airborne AVIRIS-NG and GAO campaigns.
Figure 12 gives the percentage of sources observed in each campaign and the distribution of source sectors for sources observed across those campaigns.
We observe a significantly higher false negative rate for plumes associated with Electricity Generation (EG) sources – roughly higher on training data, and nearly higher on test data – relative to the other sectors. The issues with detecting plumes from EG sources are the result of two issues. First, all of the plumes we observed from EG sources occur in urban areas with diverse surface materials and substantial albedo variation. As shown in Figure 1, CMF images representing complex and diverse urban areas often contain similarly complex false enhancements, with many occurring in the relative vicinity of observed plumes. Second, EG sources only represent 1.3% of the all of the sources observed in the selected airborne campaigns. Given the complexity of the CMF enhancements where EG sources occur, it is unlikely we have a representative sample of plumes produced by EG sources to train our model to detect them robustly. Interestingly, we observe the lowest FNR of all sectors on wastewater treatment facilities (% of all observed airborne sources) which also occur in similarly diverse urban areas, but are also likely driven by the lack of representative samples. In terms of the overall quantity of false negatives per campaign, oil and natural gas (Oil&NG) sources dominate by a factor of on training plumes, and nearly on test plumes, relative to the average false negative count of the remaining source sectors. This is not particularly surprising given that Oil&NG sources represent 86% of the all of the identified sources observed in the airborne campaigns we consider in this work. A less obvious contributing factor to the high Oil&NG FNR count is that many false negatives represent Oil&NG plumes observed in the AVIRIS-NG and GAO Permian campaigns with small pixelwise footprints. Additionally, observations from the AVIRIS-NG Permian campaign are lower resolution than those from other AVIRIS-NG campaigns (7.4 vs. 3.1 GSD), and observations from the GAO Permian campaign are lower resolution (6.0 vs. 4.67 GSD) than the other GAO campaigns.
The relative impact of plume size and concentration on plume detectability vary according to the capabilities of plume detector model, spatial resolution, sensor platform, and retrieval approach. While challenges involving the detection of specific emission source types are broadly relevant, detectability tradeoffs between plume area and concentration level are sensor and data specific, and our results shown in 11 are only directly applicable to AVIRIS-NG & GAO CMF products derived using the SWIR2-only retrieval from fine-resolution observations with GSD 3-7. That said, we can reasonably expect that replicating our AVIRIS-NG and GAO-based experiments with analogous data captured by advanced airborne sensors like AVIRIS-3, or replacing the SWIR2 CMF retrievals with full-VSWIR CMF retrievals in the experiments above will improve detectability at all plume sizes and concentrations. However, the relative impact of plume concentration and area on plume detectability are less clear for the full-VSWIR retrievals derived from coarse resolution spaceborne observations captured by spaceborne platforms like EMIT and the Carbon Mapper Coalition’s Tanager-1. Despite their advanced sensing capabilities relative to analogous airborne sensors like AVIRIS-NG and GAO, subpixel averaging in coarse resolution spaceborne radiance data at average GSD of 60 inevitably reduces retrieved concentration levels relative to retrievals derived using fine resolution airborne spectrometer data. Consequently, small or low concentration plumes that are detectable in airborne spectrometer observations are often undetectable (or unidentifiable) in spaceborne observations.
7 Discussion
In this work, we present a GHG plume detection system developed with the fundamentals necessary for operationalization. A state-of-the-art deep learning plume detection model optimized for both pixelwise segmentation and instance detection stands on a strong foundation of multicampaign, quality controlled, and appropriately sampled methane plume data. Here, we discuss literature gaps we address, open issues that remain, and future research directions to address current challenges.
Background Sampling: Background samples play a crucial role in constructing and validating a robust plume detector. Collecting a representative negative class dataset (i.e. the background) is as important as collecting a representative positive class dataset (i.e. plumes), and failure to do so will result in poor generalization. For example, sampling an equal number of plume and background tiles for model development results in a biased representation of the true observed distribution, where plume pixels are a distinct minority (Růžička et al., 2023). Sampling such few background tiles also misses many of the rare false enhancements that drive most false positive detections. Therefore, a model trained and evaluated on datasets with equal quantities of the plume and background will appear to perform well, but it will generalize poorly operationally when encountering the true observed distribution and the most challenging false enhancements. We illustrate this in detail in Suppl. Section LABEL:apx:bias_sample and Figure LABEL:fig:instance_tilebias.
Our work addresses this potential issue by sampling as many background tiles from scenes as possible. As described in Section 3 and Table 5, the plume-to-background ratio of the AVIRIS-NG dataset used to train our models was less than 1:13, an unbalanced sampling that more accurately reflects the observed rarity of high concentrations, which allows our models to perform well when evaluated on entire scenes without tile sampling in Section 5.1.
Tilewise and Scenewise Capabilities: Even with proper sampling, tilewise plume detection performance metrics (Figure 8) tend to be optimistic estimates of scenewise plume detection accuracy (Figure 6). This gap between tilewise and scenewise performance is exacerbated by any undersampling of the background class as discussed previously. Since operational plume detection systems are deployed on entire scenes and not sampled tiles, it is crucial to report scenewise capabilities (as in Section 5.1) in addition to any tilewise capabilities, which remain useful for comparing model performance (as in Section 5.2) but not for practical evaluation of operational applicability.
Unambiguous Detection and Segmentation Metrics: While most prior work reports the metric for which a model was optimized, measuring the operational performance of a proposed plume detector requires measuring both instance detection and pixelwise segmentation performance as defined in Section 2.5. As Figure 6 illustrates, a model which performs poorly in pixelwise plume segmentation metrics (e.g., our tilewise-trained model) can significantly outperform comparable models in plume instance detection metrics. Recognizing that certain use cases may prioritize either segmentation (e.g. plume masking for emission rate estimation) or detection (e.g. identifying sources for review), reporting both metrics is necessary for complete operational evaluation.
Capturing detection and segmentation accuracy with respect to false positive and false negative prediction errors is also important to provide an unbiased view of a proposed model’s capabilities. “Coupled” metrics such as traditional classification accuracy can provide a misleading view of performance, especially when samples are imbalanced. In cases where a single metric is necessary for comparison, F- scores, where weighs recall relative to precision, may be appropriate in addition to precision and recall.
Replicable Data Collection & Sampling: Assessing the operational effectiveness of a ML-driven GHG plume detection approach requires detailed accounts that describe the assumptions and practices involved in data collection, tile sampling and cross-validation. Specifics describing how samples were collected and the scope and diversity of both the plume class and the background class are necessary to determine if they are sufficiently representative to construct the model and assess its effectiveness and the extent which known sources of bias might impact reproducibility on similar data, and potential replicability in broader operational settings. For instance, if a proposed approach does not specify whether sampled tiles are allowed to overlap within scenes, across distinct scenes, or between plume and background samples, we cannot assess the impact of spatiotemporal bias on the associated model’s predictions, nor can we replicate results on similar data. To-date, detailed data collection and model validation procedures have been sorely lacking in the literature; we aimed to address this by thoroughly describing data sampling strategies in Section 3, supported by Suppl. Sections LABEL:apx:datainfo and LABEL:apx:qc.
Observational Validation of WRF-LES-based Plume Detectors: A common theme among proposed plume detection approaches is to construct synthetic GHG plume observations by injecting plumes with known characteristics generated by the WRF-LES system into plume-free “background/noise” retrieval products derived from data captured by the target platform. However, all such works in the literature are limited in evaluating their WRF-LES-based detection systems on observed data captured by the target remote sensing platforms for operational deployment. Prior works that only evaluate their WRF-LES-trained models on other WRF-LES plumes (Jongaramrungruang et al., 2022; Bruno et al., 2023) are difficult to discern for operational feasibility due to the significant differences between plumes generated by WRF-LES and real-world plumes. Other prior works do discuss results of their WRF-LES-trained models on observed methane plumes (Joyce et al., 2022; Radman et al., 2023; Xu et al., 2024); however, they only report a couple dozen true positive detections given small scenes containing previously identified plumes, or perform emission rate quantification on manually detected plumes. The omission of false positive or false negative detections on a large number of instrument-observed scenes makes it difficult to determine the capabilities of these methods in an operational setting.
In order to determine the capabilities of WRF-LES trained models when applied to observational data, we reproduced the plume detection results of Jongaramrungruang et al. (2022) and compared the capabilities of the WRF-LES-trained models to AVIRIS-NG-trained models in Suppl. Section LABEL:apx:bias_distrib. The results demonstrate poor generalization, with the WRF-LES-trained models failing to perform beyond a 0.3 F1-Score in tilewise instance detection on AVIRIS-NG test datasets despite reporting a nearly 1.0 F1-Score on the WRF-LES test dataset. This demonstrates that discrepancies between WRF-LES simulations and CMF GHG retrievals pose a significant challenge to the operational deployment of models trained solely on WRF-LES plumes, despite being injected on real-world backgrounds. Such discrepancies include the lack of changes in wind speed and direction at different times and altitudes, as well as the inability to simulate non-point-source diffuse emissions, which make up a significant portion of real world emissions and the AVIRIS-NG dataset.
Avoiding Spatiotemporal Bias and Data Leakage: CNN-based plume detectors are prone to memorizing, or “overfitting” to, site-specific features that occur in spatially overlapping observations of the same site, or from overlapping image tiles extracted from the same retrieval image scene. When overlapping images are present in both training and test sets, data leakage occurs. Validation procedures corrupted by data leakage generally produce highly optimistic validation metrics, as the model is able to “cheat” on test samples that overlap training samples with spatially correlated features. For example, Schuit et al. (2023) reports precision and recall scores over 0.97 using a CNN to distinguish tiles of labeled plume and non-plume tiles from TROPOMI XCH4 products between 2018-2020. However, the plume samples are derived from observations of only 60 regions with persistent emission sources, making it likely that many of the plumes represent repeat observations of the same emission sources. When tiles with 50% overlap are sampled from each TROPOMI image, then randomly split into training and test sets without stratification, data leakage inevitably occurs, likely accounting for the abnormally high precision and recall detection performances. When applied to TROPOMI XCH4 products from 2021, this work reports a more realistic precision score of 0.61 for confident plume detections and 0.76 for both confident and ambiguous detections. However, these results are reported after an undefined filtering process and calculated from a post hoc analysis of detection results instead of a previously labeled dataset, therefore omitting false negative analysis.
Our work makes a significant effort to control for spatiotemporal bias and prevent data leakage in our tile sampling and cross validation procedures (described in Section 3) by ensuring that spatially overlapping observations are assigned to either the training or the test dataset, but not both. As a result, a plume nor its source location (whether emitting or not) will never appear in both the training and test datasets. We also demonstrate the potential impacts of data leakage in Suppl. Section LABEL:apx:bias_stratify by inducing data leakage with a random non-stratified train-test split and showing its results. One such experiment with data from the Four Corners campaign showed that failing to prevent data leakage resulted in a misleading inflation of the model F1-Score by 83%.
Spatiotemporal bias and data leakage also likely contribute to the high overall detection accuracies reported in WRF-LES-based results, as most prior works inject multiple versions of WRF-LES generated plumes into the same observed “background” retrieval image, or worse, repeatedly inject a single WRF-LES plume instance into multiple distinct backgrounds. Allowing these samples to cross training and test boundaries will introduce severe validation bias, as the model will be tested on plumes and backgrounds identical to those in the training set. The former is especially common in prior works, with many training datasets composed of many WRF-LES simulated plumes injected into a smaller collection of background images (Joyce et al., 2022; Radman et al., 2023; Bruno et al., 2023; Jongaramrungruang et al., 2022). If the same plume-background images appear in both the training and test datasets, it increases the likelihood that the CNN learns to make plume detections based on background surface features seen during training (e.g. roads, infrastructure, geographic features) instead of the plume itself (by “overfitting”). Deliberate care must be taken to ensure background tiles do not leak between the training and test datasets.
7.1 Open Issues
One challenge we encountered in comparing our results to published work is the use of nonstandard retrievals as input to plume detector models. For instance, Růžička et al. (2023), use mag1c (Foote et al., 2020) retrievals, while Kumar et al. (2023) used newly contributed linear regression model. As a result, we cannot easily attribute differences in detection performance to the plume detection model as opposed to deviations in products derived by the respective retrieval approaches. While alternative retrieval methods likely address some limitations of the CMF retrieval approach, the relative benefits and costs involved in replacing the CMF with an alternative retrieval have only been established a handful of curated scenes.
Since false enhancements are the primary driver of prediction errors, retrieval methods that suppress or eliminate false enhancements simplify plume detection. For instance, while EMIT uses exactly the same matched-filter-based retrieval algorithm as the airborne CMF retrievals we consider in this work, the EMIT retrievals exploit the full VSWIR wavelength range of the sensor rather than just the SWIR2 absorption used in the airborne retrievals. EMIT also leverage dynamically-generated template spectra that account for scenewise characteristics Foote et al. (2021). These updates significantly reduce the number of false enhancements from surface confusers observed in typical EMIT scenes, and retroactive application to existing airborne observations has shown similar results. Additionally, recent work by Fahlen et al. (2024) has the potential to suppress false enhancements driven by varying albedo or illumination conditions.
Some classes of false enhancements cannot be rejected using models driven by the CMF alone. Augmenting CMF-driven CNN with diagnostic input bands can suppress some classes of currently unidentifiable enhancements, as shown by Satish et al. (2023).
7.2 Future Work
Given our objective of constructing plume detector models that perform as well as domain experts, controlled user studies that assess which plumes are consistently identified by multiple domain experts and the variability among their provided plume labels would provide valuable insights. Prior works have demonstrated the utility of GHG plumes manually identified by domain experts, and the results of such studies would (i) inform operational accuracy requirements for deployed models, (ii) provide a basis to assess relative tradeoffs between plume detection and pixelwise segmentation objectives, and (iii) connect results from automated, statistical plume detection approaches with prior analysis estimating the probability of plume detection, which are currently driven by manual plume identification procedures.
While advances in sensing capabilities and retrieval algorithms will provide commensurate improvements in plume detection capabilities, much work remains to address the scope and diversity of global observations captured by spaceborne platforms like EMIT and the Carbon Mapper Coalition’s Tanager-1. The improved plume detection and segmentation accuracy on the EMIT data relative to the AVIRIS-NG results (Figure 6) provide some supporting evidence that detection capabilities improve in tandem with advanced hardware and retrieval methods. Indeed, many of the problematic false enhancements we encountered in early airborne imaging campaigns were driven by limitations of the AVIRIS-NG SWIR2-only CMF retrieval approach. These improvements also demonstrate that the quality control procedures applied to the publicly released EMIT products are effective.
Data triage routines will also play a crucial role in a fully operational plume detection system. Rather than applying a computationally expensive plume detector en-masse to all data products, efficient data triage routines can optimize compute resources by selectively applying the detector to scenes with adequate SnR. For instance, scenewise summary statistics are often sufficient to detect severe data quality issues such as high background enhancement scenes covered by false enhancements (e.g. Figure 10 B.2). Cross-referencing high concentration enhancements with relevant external data products such as derived mineralogical maps, cloud masks, and geoprocessing metadata (e.g., pixelwise sensor-to-ground path lengths) can disambiguate otherwise unidentifiable false enhancements. Targeted data triage routines that leverage relevant external data sources can effectively reduce false positive detections driven by limitations of the GHG retrieval approach.
8 Conclusions
Automating GHG plume detection from remotely sensed imagery has proven to be a deceptively difficult machine learning problem requiring high-quality, interpretable labeled image retrieval products that capture diverse GHG emission sources and their surrounding background enhancements from distinct regions under varying imaging conditions, along with systematic, rigorous validation procedures aligned with operational science objectives. Interdisciplinary expertise is essential to define protocols that unambiguously characterize generalization performance of a candidate plume detection system, and to identify and account for hidden model and data-driven biases that impact generalization capabilities.
This work demonstrated the potential of ML-driven plume detection as a viable route to automate GHG plume detection in operational settings. Provided representative plume and background image retrieval products derived from data captured by current imaging spectrometers, we showed that established ML models equipped with rigorous optimization and validation procedures lead to results that are both reproducible and replicable on observations with analogous characteristics.
While previously published work claimed promising proof-of-concept results using ML-driven approaches for plume detection, detailed review revealed a variety of methodological issues that lead to significant biases and overly optimistic results for real-world deployments. Furthermore, due to the absence of crucial details regarding data collection, model optimization, and/or validation procedures, the majority of published results are not reproducible experimentally. Prior efforts all lack rigorous validation procedures to ensure that trained models generalize on independent and identically distributed observational data that generally lead to optimistic results driven by data and model-driven biases. Consequently, even in cases where the published results are feasible to reproduce, successfully replicating published accuracies on distinct but analogous data is unlikely. Here, we take steps to address these gaps in the literature, with the goal of pushing towards operational deployments of automated plume detection models.
Data and Code Availability
AVIRIS-NG and EMIT datasets as described, both for tilewise model training and scenewise evaluation, are available at https://ml.jpl.nasa.gov/datasets/opsghgml/. GAO datasets are not available for public access due to data controls, and only evaluation results are presented as reference. Dataset sampling and preprocessing code, as well as model training, inference, and evaluation code, is open source and available at https://github.com/JPLMLIA/operational-ghg.
Acknowledgments
The Earth surface Mineral dust source InvesTigation (EMIT) plume delineations were performed using a tool built with the Jet Propulsion Laboratory Multi-Mission Geographical Information System. Carbon Mapper acknowledges the generous support of its philanthropic donors. The High Performance Computing resources used in this investigation were provided by funding from the JPL Information and Technology Solutions Directorate. EMIT is supported by the National Aeronautics and Space Administration Earth Venture Instrument program, under the Earth Science Division of the Science Mission Directorate. The research was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). © 2025. California Institute of Technology. Government sponsorship acknowledged.
Appendix A Data Table Column Definitions
1.
Sensor Platform: imaging spectrometer sensor platform used to collect survey data set.
2.
Survey Data Set: survey / data set identifier.
3.
Survey Regions: regions indicated by US state name or ’global’ measured in data set.
4.
Survey Start/End (Date): begin / end dates of image acquisitions captured in data set.
5.
Unique Scenes (#): Total number of labeled CMF scenes in data set.
6.
GSD (): Average Ground Sampling Distance (GSD) in of all scenes in data set.
7.
Length (km): Average length in km of all scenes in data set.
8.
Unique Plumes (#): Total number of manually identified methane plume candidates in labeled scenes in data set.
9.
Plume Tiles (#): Total count of plume tiles in data set.
10.
Bg Tiles (#): Total count of background tiles in data set.
11.
Quality Control (# Reviewers): Quality control level and (number of reviewers / scene).
12.
Train:Test (Val) Scenes (#): Total training and test scenes used in tilewise optimization and (total test scenes used in scenewise validation).
Appendix B Airborne Plume Source Sectors
The distribution of source sectors for plumes observed in airborne imaging campaigns are provided in Figure 14.